OpenStudy (anonymous):

7. What is the slope of the line that passes through the pair of points (2, 5) and (8, 3)? A. 1/3 B. -1/3 C. 3 D. -3

OpenStudy (anonymous):

8. What is the slope of the line that passes through the pair of points (-7/2, -3) and (-5, 5/2)? A. 6/22 B. -6/22 C. 22/6 D. -22/6 Thanks!

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

pooja195 (pooja195):

1) \[\LARGE \frac{3-5}{8-2}=slope\]

OpenStudy (anonymous):

D and C??

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Nope

pooja195 (pooja195):

3-5=?

OpenStudy (anonymous):

-2

pooja195 (pooja195):

8-2=?